OpenStudy (anonymous):

How do you multiply (2-3i)^2

OpenStudy (solomonzelman):

Wait, you are doing complex and imaginary numbers right now, or just algebra or pre-algebra?

OpenStudy (anonymous):

i is the imaginary number, its algebra 2.

OpenStudy (solomonzelman):

Okay, good...

OpenStudy (solomonzelman):

there is a rule: \(\large\color{black}{(a-b)^2=a^2-2ab+b^2 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) you know that \(\large\color{black}{i=\sqrt{-1} }\)\(\LARGE\color{white}{ \rm \left| \right| }\) and just like \(\large\color{black}{\sqrt{a} \times \sqrt{a}=a}\) so is,\(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{i \times i}\) (which is \(\large\color{black}{\sqrt{-1} \times \sqrt{-1}}\)) is\(\LARGE\color{white}{ \rm \left| \right| }\) equal to \(\large\color{black}{-1.}\)\(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{\sqrt{a} \times \sqrt{a}=a}\)\(\LARGE\color{white}{ \rm \left| \right| }\)

OpenStudy (solomonzelman):

the last line shouldn't be there, read everything without the very last equation, please.

OpenStudy (solomonzelman):

\(\large\color{black}{(2-3i)^2=(2)^2-2(2)(3i)+(3i)^2}\)

OpenStudy (solomonzelman):

if you have any questions after you are done reading, please ask.